What is the probability that an adult’s weight is in between 55-65kg?


1
Expert's answer
2021-02-28T06:58:36-0500

The weights of a normally distributed group of adults participating in a fitness test has an average of 60kg and a standard deviation of 8kg.

a. What is the probability that an adult’s weight is in between 55-65kg?


Here, Mean("\mu" )= 60

Standard Deviation("\sigma" )= 8

Using the formula "z=\dfrac{x-\mu}{\sigma}"

"z_1=\dfrac{55-60}{8}=-0.625"


and "z_2= \dfrac{65-60}{8}=0.625"




Hence, P(-0.625<z<0.625)= 0.4714